OpenStudy (anonymous):

differentiate y= (x^2+4x+3)/x^(1/2)

OpenStudy (anonymous):

best bet is to divide each term by \[x^{\frac{1}{2}}\] by subtracting the exponents, then use the power rule for each term

OpenStudy (anonymous):

\[y=x^{\frac{3}{2}}+4x^{\frac{1}{2}}+3x^{-\frac{1}{2}}\] and then power rule three times

OpenStudy (anonymous):

how did you get the 3x^(-1/2) if the 3 does not have an X?? im confused

OpenStudy (anonymous):

y= (x^2+4x+3)/x^(1/2) y=x^2/x^(1/2)+4x/x^(1/2)+3/x^(1/2) y=x^(2-1/2)+4x^(1-1/2)+3x^-1/2 y=x^3/2+4x^1/2+3x-1/2 now when u diff y' y'=3/2x^1/2+4*1/2*x^-1/2-3/2x^-3/2 y'=3/2x^1/2+2x^(-1/2)-3/2x^-3/2